OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m= -5, (4,9)

OpenStudy (godlovesme):

Use the point slope formula R u familiar with that :)

OpenStudy (anonymous):

Yes its just been a while

OpenStudy (godlovesme):

K y-y1= m(x-x1) M is ur slope Y1 ur given y coordinate X1 given x coordinate Can u plug those to the eqn

OpenStudy (anonymous):

yeah I know that one

OpenStudy (anonymous):

just have to use point slope formula

OpenStudy (solomonzelman):

sheesh I suck, better move to english for now, will confuse no more

OpenStudy (anonymous):

I will try

OpenStudy (godlovesme):

(4,9) slope=-5 y-coordinate=9 x-coordinate=4 y-y1=m(x-x1)

OpenStudy (anonymous):

what answer did you get so I can compare

OpenStudy (godlovesme):

y-9=-5(x-4) y-9=-5x+20 y=-5x+29 u?

OpenStudy (anonymous):

y=-5x-49

OpenStudy (godlovesme):

how did u get -49

OpenStudy (anonymous):

y -4 = -5x -45

OpenStudy (godlovesme):

how did u set up ur equation

OpenStudy (anonymous):

y - 4 -5(x + 9)

OpenStudy (godlovesme):

k distribute -5 and dont forget =

OpenStudy (anonymous):

Alright

OpenStudy (godlovesme):

im sorry u r right @Bigtim10

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Thanks for your help though

OpenStudy (godlovesme):

wait

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

So, may I recommend @Bigtim10, state point-slope form to begin

OpenStudy (godlovesme):

(4,9) slope=-5 y-9= -5(x-4) thats how it was supposed to be like =0

OpenStudy (anonymous):

no

OpenStudy (fibonaccichick666):

who is the no directed at?

OpenStudy (godlovesme):



OpenStudy (anonymous):

y - 4 = -5(x - 9)

OpenStudy (godlovesme):

y-y1= m(x-x1) y-9= -5(x-4) @Bigtim10

OpenStudy (fibonaccichick666):

so, that is why you should type out the formula first bigtim

OpenStudy (fibonaccichick666):

you have a small error, \[y-y_0=m(x-x_0)\]

OpenStudy (fibonaccichick666):

you switched the x-not and y-not

OpenStudy (anonymous):

I remember from when I did early in the class

OpenStudy (fibonaccichick666):

but I think you can do this, in fact I'm sure you can, so give it a try for us. Show us until you get stuck, then we can give you hints

OpenStudy (godlovesme):

@Bigtim10 sorry for the confusion >.<

OpenStudy (anonymous):

no problem. Thanks

OpenStudy (godlovesme):

anytime :)